OpenStudy (anonymous):

Why is (6, –2) not a solution to the system of equations given below? 2x-3y=18 5x+3y=3 A. It satisfies the top equation but not the bottom equation. B. It satisfies the bottom equation but not the top equation. C. It doesn’t satisfy either the top or the bottom equations. D. It satisfies both the top and the bottom equations.

OpenStudy (jdoe0001):

(6, -2) x y plug them in, see if they work

OpenStudy (whpalmer4):

Plug in the values and see what happens. Note that all of the equations must work, not just some of them!

OpenStudy (anonymous):

yes they do

OpenStudy (anonymous):

2x - 3y = 18 (1) 5x + 3y = 3 (2) Substitute x = 6 and y = -2 into the 2 equations The top equation (1) is satisfied with (6, -2) The bottom equation (2) isn't. Then, the answer is......